# Non-autoregressive Generative Models for Reranking Recommendation

### v1

@@ 4.4. Sequence Evaluator
The sequence evaluator model is designed to estimate the overall utility of a given sequence, as illustrated in fig. 2. The generated sequence from the generator undergoes encoding through self-attention and a feed-forward layer to obtain contextual information. The positional representation passes the linear projection layer to predict the score for a specific target. The overall utility is calculated as the weighted sum of scores at each position.
@@ 5. Experiments
We conduct extensive experiments in both offline and online environments to demonstrate the effectiveness of NAR4Rec. Three questions are investigated in the experiments:
  • How does NAR4Rec perform in comparison with state-of-the-art performance in terms of recommendation quality?
  • How do different settings of NAR4Rec affect the performance?
  • How does NAR4Rec perform in real-life recommendation systems?
@@ 5.1. Experiment details
Dataset: To ensure dataset consistency and align with the problem setup, we emphasize the inclusion of real item sequences displayed in complete sessions, rather than manually constructed lists. Consequently, we have collected two datasets specifically for experiments.
  • Avito Dataset: The Avito dataset is a publicly available collection of user search logs. Each sample represents a search page containing multiple ads, forming a really impressive list with feedback. The dataset comprises over 53 million lists involving 1.3 million users and 36 million ads. The training phase utilizes impressions from the first 21 days, while the testing phase utilizes impressions from the last 7 days. Each list in the dataset has a length of 5.
  • VideoRerank Dataset: The VideoRerank dataset is derived from Kuaishou, a widely used short-video application with a user base of over 300 million daily active users. Each sample in the dataset represents an actual sequence of items displayed to a specific user during a complete session, ensuring dataset consistency and problem alignment. The dataset consists of a total of 82,230,788 users, 26,835,337 items, and 1,811,625,438 requests. Each request includes 6 items and 60 candidates, about the ranking process.
Baselines We conduct a comparative analysis of our NAR4Rec approach with five widely adopted reranking methods, namely DNN, DCN, PRM, PIER, and Edge-Rerank, which serve as our baselines. Here is a brief overview of these baseline methods:
  • DNN(Covington et al., 2016): DNN is a fundamental method for click-through rate (CTR) prediction that utilizes a multi-layer perceptron (MLP) to capture high-order feature interactions.
@@
  • PRM(Pei et al., 2019): PRM enhances an initial list by leveraging the self-attention mechanism to account for the mutual influence among items.
  • Edge-Rerank(Gong et al., 2022): Edge-Rerank combines an on-device ranking model with an adaptive beam search approach to generate context-aware re-ranking outcomes.
  • PIER(Shi et al., 2023): PIER treats the target ad and all slots as a query and employs external items as key-value pairs to model externalities in different exposure situations.
Metrics In our experiments, we employ commonly used metrics to evaluate the performance of different methods on the Avito dataset and the VideoRerank dataset. Specifically, we utilize AUC and LogLoss as evaluation metrics for the Avito dataset, while for the VideoRerank dataset, which contains both candidates and exposed item lists, we employ Recall@6, Recall@10, and LogLoss as evaluation metrics.
5.2. Offline experiments
In this section, we conducted offline experiments to evaluate the effectiveness of our method on two datasets.
Hyperparameters In our offline experiments with the Avito dataset, we use the following hyperparameters: a learning rate of 0.001, a batch size of 1024, and the Adam optimizer. For the VideoRerank dataset, we maintain a learning rate of 0.001, a batch size of 256, and utilize the Adam optimizer as well.
5.2.1. Performance comparison(RQ1)
The experimental results on the Avito dataset are presented in table 1, following the approach proposed by (Shi et al., 2023). AUC and LogLoss are the evaluation metrics used in this study. Our method achieves the best performance among the baseline models in both AUC and LogLoss. PRM outperforms DNN and DCN by effectively capturing the mutual influence between items. Additionally, Edge-Rerank surpasses DNN and DCN by integrating an on-device ranking model with an adaptive beam search technique. PIER demonstrates superiority over other baselines by considering the interaction per category. Notably, our proposed method exhibits the highest improvement with a significant increase of 0.0125 in the AUC metric compared to other baseline models.
@@ 5.2.1. Performance comparison(RQ1)
AUC LogLoss
DNN 0.6614 0.0598
DCN 0.6623 0.0598
PRM 0.6881 0.0594
Edge-rerank 0.6953 0.0574
PIER 0.7109 0.0409
NAR4Rec 0.7234 0.0384
5.2.2. Analysis of NAR4Rec(RQ2)
As shown in fig. 3 and fig. 4, by conducting a series of experiments on NAR4Rec, we demonstrate that our experimental results exhibit insensitivity to variations in the learning rate and batch size.
Figure 3. The performances of NAR4Rec on Avito Dataset with different learning rate.
Figure 4. The performances of NAR4Rec on Avito Dataset with different batch size.
Table 2 shows the results of our offline experiments on VideoRerank. The evaluation metrics used in this experiment include Recall@6, Recall@10, and Loss. Our method achieves superior results on all metrics compared to other baseline models as well.
Table 2. Comparison between Nar4Rec and other baselines on VideoRerank
@@ 5.2.2. Analysis of NAR4Rec(RQ2)
5.3. Online experiments(RQ3)
We conduct online A/B experiments on the KuaiShou App.
@@ 5.3.1. Experiments setup
In online A/B experiments, we evenly divided the traffic of the entire app into ten buckets. The online baseline is Edge-rerank(Gong et al., 2022), with 20% of the traffic assigned to NAR4Rec, while the remaining traffic continues with Edge-rerank. We set two conditions for NAR4Rec: first, we expect the users to watch more videos and stay more time on each video; second, we expect positive feedbacks from users.
@@ 5.3.2. Experimental Results
The experiments have been launched on the system for ten days, and the result is listed in table 3. NAR4Rec outperforms Edge-rerank(Gong et al., 2022) by a large margin.
5.3.3. Ablation study
Given the absence of an evaluation metric for the generated sequence, we conducted online experiments to demonstrate the effectiveness of NAR4Rec. The removal of the contrastive search resulted in a decrease in diversity by +0.847%. Furthermore, training on all exposed item lists led to a significant drop in watch time by -1.0% (considering a drop of around 0.2% as significant in our system).
Views Likes Follow Long Views Complete Views
@@ 5.3.3. Ablation study
Table 3. Online experiments results. All values are the relative improvements of NAR4Rec. For online A/B test in Kuaishou, the improvements of over 0.5% in position interactions(like, follow) and 0.2% in views are very significant.
@@ 6. Conclusion
In this paper, we have presented an overview of the current formulation and challenges associated with reranking in recommendation systems. Though non-autoregressive generation has been explored in natural language processing, the conventional techniques are not directly applicable to recommendation systems. We tackle the challenges encountered by recommendation to improve the convergence and performance of non-autoregressive models and make the first attempt to integrate non-autoregressive models into reranking in real-time recommender systems. Extensive online and offline experiments have demonstrated the efficacy of NAR4Rec as a versatile framework for generating sequences with enhanced utility. Moving forward, our future work will focus on refining the modeling of sequence utility to further enhance the capabilities of NAR4Rec.

### v3

@@ 4.4. Sequence Evaluator
The sequence evaluator model is designed to estimate the overall utility of a given sequence, as illustrated in fig. 2. The generated sequence from the generator is first encoded using a self-attention and a feed-forward layer to capture contextual information. The hidden representation then passes through the linear projection layer to predict the score for a specific target. The overall utility is calculated as the weighted sum of itemwise scores. Ultimately, the sequence with the highest overall utility is chosen for delivery to the users.
@@ 5. Experiments
In this section, we conduct extensive offline experiments and online A/B tests to demonstrate the effectiveness of NAR4Rec. We first describe our experiment setup and baselines in section 5.1. For offline experiments in section 5.2, we compare NAR4Rec with existing baselines on both performance and training and inference time. Then we alternate the hyper-parameter to analyse the hyper-parameter sensitivity of NAR4Rec. To further show the effectiveness of NAR4Rec in real-time recommendation system, we conduct online A/B tests to ablate our proposed methods in section 5.3.
@@ 5.1. Experiment details
Table 1. Statistics of the datasets.
Dataset #Requests #Users #Ads
Avito 53,562,269 1,324,103 23,562,269
Meituan 230,525,531 3,201,922 98,525,531
Dataset: To evaluate reranking recommendation, we expect that each sample of the dataset is an exposed sequence to users rather than a manually constructed sequence. For public dataset, we choose Avito dataset. For industrial dataset, we use real-world data collected from Kuaishou short-video platform. The detailed introduction is given in table 1.
  • Avito [^2] https://www.kaggle.com/c/avito-context-ad-clicks/data: The Avito dataset is a publicly available collection of user search logs from avito.ru. The dataset comprises over 53 million lists with 1.3 million users and 36 million ads. Each sample corresponds to a search page with multiple ads. The user search logs from first 21 days are used as training set and the search logs from the last 7 days are used as test set. The sequence length in Avito is 5.
  • Kuaishou: The Kuaishou dataset is derived from Kuaishou, a widely used short-video application with a user base of over 300 million daily active users. Each sample in the dataset represents an actual user request log, which contains user information(e.g. user id, age, gender), candidates items and user interaction to exposed items. The dataset consists of a total of 82,230,788 users, 26,835,337 items, and 1,811,625,438 requests. Each request contains 6 items in the exposed item sequence and 60 candidates from ranking.
5.2. Offline experiments
Baselines We compare the proposed NAR4Rec with 6 state-of-the-art reranking methods. We select DNN, DCN as pointwise baselines, PRM as one-stage listwise baselines, Edge-Rerank, PIER, Seq2slate as two-stage baselines. Crucially, Seq2slate is a autoregressive generative model. A brief overview of these baseline methods is as follows:
  • DNN(Covington et al., 2016): DNN is a basic method for click-through rate prediction, which applies a multi-layer perception to learn feature interaction.
@@ 5.2. Offline experiments
  • PRM(Pei et al., 2019): PRM models the mutual correlation among items by leveraging the self-attention mechanism and then rank the items by the estimated scores to generate the item sequence.
  • Edge-Rerank(Gong et al., 2022): Edge-Rerank generates the context-aware sequence with adaptive beam search on estimate scores.
  • Seq2Slate(Bello et al., 2018): Seq2Slate leverages pointer networks, which are seq2seq models with an attention mechanism to predict the next item given the items already selected.
  • PIER(Shi et al., 2023): PIER applies hashing algorithm to slect top-k candidates from the full permutation based on user interests. Then the generator and evaluator are jointly trained to generate better permutations.
Metrics As there is not common sequence generation metrics for recommendation, we follow previous work(Shi et al., 2023; Lin et al., 2023) and evaluate these models using three commonly adopted metrics: AUC, LogLoss, and NDCG on Avito dataset. For Avito dataset, where $n$ and $m$ are equal(5), the task is to predict the itemwise click-through rate given a listwise input. For Kuaishou dataset, where $n$ and $m$ are $60$ and $6$ respectively, we employ Recall@6, Recall@10, and LogLoss as evaluation metrics. The task for Kuaishou dataset is to predict whether an item is chosen to be one of the exposed $6$ items.
Hyperparameters For Avito dataset, the learning rate is 10-3, the optimizer is Adam and the batch size is 1024. For Kuaishou dataset, the learning rate and optimizer is the same as Avito, but the batch size is 256.
5.2.1. Performance comparison
Here we show the results of our proposed method NAR4Rec. As can be seen in in table 2 and table 3, NAR4Rec outperforms 5 baslines including recent strong reranking methods(Pei et al., 2019; Bello et al., 2018; Shi et al., 2023). PRM outperforms DNN and DCN by effectively capturing the mutual influence between items. Additionally, Edge-rerank surpass DNN and DCN with an adative beam search with previous item information. PIER demonstrates superiority over other baselines by the interaction per category. Notably, our proposed method exhibits the highest improvement with a significant increase of 0.0125 in the AUC metric compared to other baseline models. table 3 shows the results of our offline experiments on Kuaishou. The evaluation metrics used in this experiment include Recall@6, Recall@10, and Loss. Our method achieves superior results on all metrics compared to other baseline models as well.
@@ 5.2.1. Performance comparison
AUC LogLoss NDCG
DNN 0.6614 0.0598 0.6920
DCN 0.6623 0.0598 0.7004
PRM 0.6881 0.0594 0.7380
Edge-rerank 0.6953 0.0574 0.7203
PIER 0.7109 0.0409 0.7401
Seq2Slate 0.7034 0.0486 0.7225
NAR4Rec 0.7234 0.0384 0.7409
Table 3. Comparison between Nar4Rec and other baselines on KuaiShou.
@@
5.2.2. Training and Inference Time comparison
Given that NAR4Rec is closely related to autoregressive models, we conduct a comparison with autoregressive models Seq2Slate. We compare the training and inference time on the Avito dataset between Seq2slate and NAR4Rec. We also give training and inference time for generators in other baselines in table 4. Since Seq2slate utilizes recurrent neural networks as its backbone network, both training and inference processes adopt an autoregressive manner. The inference speedup of NAR4Rec over Seq2slate is almost the same as training. NAR4Rec only requires 58 minutes to complete the training while Seq2Slate requires 283 minutes. Such a significant reduction in training time (i.e. approximately 5$\times$ speedup) highlights the computational efficiency of NAR4Rec. The autoregressive model represented by Seq2Slate generates target sequences item by item, while our Non-autoregressive NAR4Rec generates all items at once. So when generating a sequence with length 5, NAR4Rec shows approximately 5$\times$ speedup.
Table 4. The training and inference time comparison between NAR4Rec and other baselines on Avito. All experiments are conducted on Tesla T4 16G GPU and the batch size is set to 1024. The training and inference time is calculated by averaging the result over 100 steps.
Training Time Inference Time
DNN 0.102s 0.034s
DCN 0.106s 0.035s
PRM 0.109s 0.036s
Edge-rerank 0.105s 0.035s
PIER 0.160s 0.053s
Seq2Slate 0.558s 0.186s
NAR4Rec 0.112s 0.037s
5.2.3. Hyper-parameter Analysis of NAR4Rec
We further analyze the hyper-parameter sensitivity on NAR4Rec. Here, we conduct a series of experiments on NAR4Rec and PIER. As shown in fig. 3, we demonstrate that our experimental results exhibit insensitivity to variations in the learning rate, batch size, and epoch.
Figure 3. The comparison between NAR4Rec and PIER on Avito with different learning rate, batch size and epoch.
Then, we analyze the impact of weight $\alpha$ and margin $\rho$ in contrastive loss and the impact of penalty parameter $\alpha$ in contrastive decoding. Figure 4 shows the results of our experiments. We change $\omega$ while fixing $\rho$=0.5, and change $\rho$ while fixing $\omega$=0.01 in contrastive loss. When changing $\alpha$ in contrastive decoding, we set $\rho$=0.5 and $\omega$=0.01 as the default parameters.
Figure 4. The effect of parameter weight $\omega$ and margin $\rho$ in contrastive loss and the $\alpha$ in contrastive decoding.
5.3. Online experiments
Text sequence generation often is evaluated by human labeling. In recommendation, we resort to online A/B experiments to obtain the feedback from users to demonstrate our effectiveness.
@@ 5.3.1. Experiments setup
In online A/B experiments, we evenly divide the traffic of the entire app into ten buckets. The online baseline is Edge-rerank(Gong et al., 2022), with 20% of the traffic assigned to NAR4Rec, while the remaining traffic is assigned to Edge-rerank.
@@ 5.3.2. Experimental Results
The experiments have been launched on the system for ten days, and the result is listed in table 5. NAR4Rec outperforms Edge-rerank(Gong et al., 2022) by a large margin. NAR4Rec shows users watch more(i.e a higher Views) videos, spend more time on each video(i.e. more Long Views and Complete Views) and a more positive user feedback(i.e. a improvement on like, follows) over (Gong et al., 2022).
Views Likes Follows Long Views Complete Views
@@
Table 5. Online experiments results. All values are the relative improvements of NAR4Rec. For the online A/B test in Kuaishou, the improvements of over 0.5% in positive interactions(like, follow) and 0.2% in views are very significant.
5.3.3. Ablation Study on Unlikelihood Training
To show the effectiveness of unlikelihood training, we compare vanilla training on all exposed sequences with unlikelihood training. Unlikelihood shows more Views and a longer Watch Time.
Views Watch Time
Vanilla training -0.370%* -0.277%*
Table 6. Online experiments on training methods. All values are the relative changes over Unlikelihood Training. * indicates that the metrics is statistically significant.
5.3.4. Ablation Study on Contrastive Decoding
Here we compare the common decoding algorithm in text sequence and a diversity-based algorithm(i.e. Deep DPP)with contrastive decoding. Those decoding algorithms shows a significant drop in View and Watch time, suggesting a poorer user feedback.
Views Watch Time
Deep DPP -0.363%* -0.361%*
Beam Search -0.327%* -0.214%*
Greedy Search -0.216%* -0.178%*
Top-k Sampling -0.254%* -0.131%
Table 7. Online experiments on decoding methods. All values are the relative changes over the Contrastive Decoding. * indicates that the metrics is statistically significant.
@@ 6. Conclusion
In this paper, we provide an overview of the current formulation and challenges associated with reranking in recommendation systems. Although non-autoregressive generation has been explored in natural language processing, conventional techniques are not directly applicable to recommendation systems. We tackle the challenges in recommendations to improve the convergence and performance of non-autoregressive models and make the first attempt to integrate non-autoregressive models into reranking in real-time recommender systems. Extensive online and offline A/B experiments have demonstrated the effectiveness and efficiency of NAR4Rec as a versatile framework for generating sequences with enhanced utility. Moving forward, our future work will focus on refining the modeling of sequence utility to further enhance the capabilities of NAR4Rec.